OpenStudy (thomas5267):

Prove that \[x^{25}\equiv x\pmod {13}\].

ganeshie8 (ganeshie8):

\[x^{25}\equiv x\cdot(x^{12})^2 \equiv x\cdot(1)^2 \equiv x \pmod{13}\]

OpenStudy (thomas5267):

How do you know x is not divisible by 13?

ganeshie8 (ganeshie8):

Oh thats a mistake actually, good catch :)

ganeshie8 (ganeshie8):

Above line works when \(13\nmid x\) Just add this for the case \(13 | x\) : \(x^{25}\equiv 0 \equiv x \pmod{13} \)

OpenStudy (thomas5267):

...... I don't know what to say really......

ganeshie8 (ganeshie8):

when 13 | x, would you agree \(x\equiv 0 \pmod {13}\) ?

ganeshie8 (ganeshie8):

if you think about them, they both are equivalent statemetns

OpenStudy (thomas5267):

Let p be a prime and \(\gcd(a,p)=1\). Use Fermat's little theorem to show that \(x\equiv a^{p-2}b\pmod p\) is a solution to \(ax\equiv b \pmod p)\)

OpenStudy (thomas5267):

Am I suppose to say by Fermat's little theorem, \(a^{p-1}\equiv 1 \pmod p\) so \(a^{p-1}b\equiv b \pmod p\).

ganeshie8 (ganeshie8):

thats it! replace \(b\) by \(a^{p-1} b\) in the given congruence : \[\large ax \equiv b \pmod{p}\] \[\large ax \equiv a^{p-1}b \pmod{p}\] divide \(a\) both sides

ganeshie8 (ganeshie8):

you're allowed to divide \(a\) both sides because \(\gcd(a, p)=1\)

OpenStudy (thomas5267):

Okay. Prove that, if \(p\geq3\) is a prime then \(\sum_{k=1}^{p-1}k^p\equiv0\pmod p\).

ganeshie8 (ganeshie8):

Easy... use Fermat's little theorem for each term and add up

OpenStudy (thomas5267):

WHAT!?!?!?!?!?

ganeshie8 (ganeshie8):

since \(1\le k\le p-1\), we have \(\gcd(k, p)=1\) so \(k^p \equiv k \pmod{p}\) : \[\sum_{k=1}^{p-1}k^p\equiv \sum_{k=1}^{p-1}k \pmod p \]

OpenStudy (thomas5267):

I am certainly thinking too complicated. I was thinking how to factorise the sum and prove that the factorisation is divisible by p lol.

ganeshie8 (ganeshie8):

Just believe that almost all the homework proofs are simple and short :)

OpenStudy (thomas5267):

It is IB past paper......

OpenStudy (thomas5267):

The difficulty is surreal yet I can't solve it.

OpenStudy (thomas5267):

I will close this question. I have an exam on discrete mathematics tomorrow and you saved my life.

ganeshie8 (ganeshie8):

I see you're fully ready for the exam! wish you all the best :)